OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts -4x-5y=30

OpenStudy (anonymous):

\[\frac{x}{\frac{30}{4}} + \frac{y}{-6} =1 \]

OpenStudy (anonymous):

so it makes x intercept (30/4 , 0 ) and 0,-6

OpenStudy (anonymous):

0,-6 is y intercept